a hot iron of mass 0.5 kg is dropped into a container filled with 5 kg of water.The container is made of 0.3 kg of iron initially at 25 celcius .The final equilibrium temperature is 30 celcius .Find the initial temperature of the hot iron


1
Expert's answer
2021-07-25T09:18:46-0400
"m_ic_i(T-T_e)=(m_ic_i+m_wc_w)(T_e-T_0)\\
(440)(0.5)(T-30)\\=((440)(0.3)+(4200)(5))(30-25)\\
T=510\degree C"